Analytic Geometry
13. On a street map, the coordinates of the two fire stations in a town are A(10, 63) and B(87, 30). A neighbor reports smoke coming from the kitchen of a house at C(41, 18).
a) Whic fire station is closer to this house?
Can you please help me with this question? Thanks>>

Answer by Alan3354(69443)

On a street map, the coordinates of the two fire stations in a town are A(10, 63) and B(87, 30). A neighbor reports smoke coming from the kitchen of a house at C(41, 18).
a) Whic fire station is closer to this house?
---------------------
Find the 2 distances.
d^2 = diffy^2 + diffx^2
Distance to A
d^2 = (18-63)^2 + (41-10)^2 = 2986
-------------
Distance to B
d^2 = (18-30)^2 + (41-87)^2 = 2260
--------------
B is closer. It's not necessary to take the sq root, since the distance wasn't asked for, only which is closer.